Nathan's Nerf Balls Company produce and sells nerf footballs and soccer balls. In the most recent month, the nerf footballs had sales of $19,000 and variable expenses of $10,030. Nerf soccer balls, on the other hand, had sales of $32,000 and variable expenses of $15,980. The fixed expenses of Nathan's Nerf Balls Company were $45,940. The break-even point for Nathan's entire company is closest to:

Multiple Choice

  • $45,940

  • $71,950

  • $93,715

  • $93,755
Answer #1

Total sales = 19000 + 32000 = 51000

Total Variable costs = 10030+15980 = 26010

Contribution margin ratio = Contribution margin/Sales

= (51000-26010)/51000

= 49%

Breakeven point = Fixed cost/Contribution margin ratio

= 45940/49%

= 93,755

Option D is the answer